OpenStudy (anonymous):

How many ounces of trail mix are in a bag that weighs 0.908 kilograms? Input only numeric values. (1 pound = 0.454 kg and 1 pound = 16 ounces) Please help me! i give fans and medals! Thank you!!!!

OpenStudy (anonymous):

@texaschic101 help me! (this is actually tal tal i just forgot my password lol)

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

I for got my password to I was Deadpool2

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

O.K. there is 0.0283495 kilograms in 1 ounce

OpenStudy (texaschic101):

1/0.454 = x/0.908 --- 1 lb to 0.454 kg = x lbs to .908 kg cross multiply 0.454x = 0.908 x = 0.908/0.454 x = 2 lbs 1 lb = 16 oz.....so 2 lb = 32 oz <=== your answer

OpenStudy (anonymous):

omg thank you so much you guys are lifesavers!!!!!!

OpenStudy (anonymous):

Well @texaschic101 did all the work so thank her "applause" "whistle"

OpenStudy (anonymous):

@texaschic101 can you help me with one more?

OpenStudy (anonymous):

Mark uses an app that shows him how many kilometers he has run to prepare for a marathon. The app said he ran 9.654 kilometers. He wants to post online how many miles he ran. Mark ran ___ miles. (1 mile = 1.609 kilometers)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

sorry I'm probably annoying you. You're just so helpful!

OpenStudy (texaschic101):

1/ 1.609 = x/9.654 --- 1 mile to 1.609 km = x miles to 9.654 km cross multiply 1.609x = 9.654 x = 9.654/1.609 x = 6 Mark ran 6 miles and don't worry, I am happy to help :)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (texaschic101):

of course :)